OpenStudy (anonymous):

Find the value of y so that the line passing through (9, 0) and (3, y) has a slope of -1/2. If someone could explain this a little bit better.

terenzreignz (terenzreignz):

Formula for the slope given two points \(\large (x_1 \ , \ y_1)\) and \(\large(x_2 \ , \ y_2)\) \[\Large m = \frac{y_2-y_1}{x_2-x_1}\]

terenzreignz (terenzreignz):

Compactly, \[\Large m = \frac{\Delta y}{\Delta x}\]

OpenStudy (anonymous):

what do those triangles mean

terenzreignz (terenzreignz):

They mean difference. It's just a shorter way of writing \[\Large m = \frac{y_2-y_1}{x_2-x_1}\] Those 'triangles' are the uppercase Greek letter Delta, which means difference, so in effect, \[\Large m = \frac{\Delta y}{\Delta x}\] just means 'difference of y's divided by difference of x's' If those 'triangles' look that much alien to you, then pay no heed to them, and just focus on this slope formula: \[\Large m = \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

well what do i do with the given y in the problem

terenzreignz (terenzreignz):

Well, you can consider that y to be your \(y_2\)

OpenStudy (anonymous):

(9,0) and (3,y) how do i subtract it like that

OpenStudy (anonymous):

the y i mean

terenzreignz (terenzreignz):

What are the y-values of these two points? By y-values, I mean the right-coordinates...

OpenStudy (anonymous):

0 and y

terenzreignz (terenzreignz):

And their difference is...?

OpenStudy (anonymous):

-1/2? or is it the other coordinates which would be 9 and 3?

terenzreignz (terenzreignz):

No, the difference of 0 and y. So y - 0 = y :P

terenzreignz (terenzreignz):

Their difference is y. Got that so far?

OpenStudy (anonymous):

so y is 0..?